# Sparse-DeRF: Deblurred Neural Radiance Fields from Sparse View

### v1

@@ Abstract
Recent studies construct deblurred neural radiance fields (DeRF) using dozens of blurry images, which are not practical scenarios if only a limited number of blurry images are available. This paper focuses on constructing DeRF from sparse-view for more pragmatic real-world scenarios. As observed in our experiments, establishing DeRF from sparse views proves to be a more challenging problem due to the inherent complexity arising from the simultaneous optimization of blur kernels and NeRF from sparse view. Sparse-DeRF successfully regularizes the complicated joint optimization, presenting alleviated overfitting artifacts and enhanced quality on radiance fields. The regularization consists of three key components: Surface smoothness, helps the model accurately predict the scene structure utilizing unseen and additional hidden rays derived from the blur kernel based on statistical tendencies of real-world; Modulated gradient scaling, helps the model adjust the amount of the backpropagated gradient according to the arrangements of scene objects; Perceptual distillation improves the perceptual quality by overcoming the ill-posed multi-view inconsistency of image deblurring and distilling the pre-filtered information, compensating for the lack of clean information in blurry images. We demonstrate the effectiveness of the Sparse-DeRF with extensive quantitative and qualitative experimental results by training DeRF from 2-view, 4-view, and 6-view blurry images.
@@ I Introduction
In real-world scenarios, tackling the blurry images from camera motion is regarded to be important since users often encounter degraded images when capturing photos with their own devices due to the unintentional camera movement during exposure time. To solve this problem, several NeRF studies [1, 2, 3] have attempted to construct deblurred neural radiance fields (hereafter, DeRF) from blurry images using joint optimization of internal implicit blur kernel and radiance fields, but they use dozens of blurry images to train, which is actually not practical scenarios. The assumed experimental environments, where radiance fields are trained from about 20 to 30 blurry images, seem unlikely to occur in reality Hence we delve into the practical consideration for situations where only blurry images are utilized. We reasoned that situations requiring the use of only blurry images would arise when the available images for reconstructing the desired 3D space are both very limited and blurry. Following this rationale, we propose a novel pragmatic scenario for radiance fields from blurry images that establish the DeRF from sparse view settings. Specifically, we set the 2-view, 4-view, and 6-view settings based on our consideration of the practical applications of research on generating radiance fields from blurry images.
Actually, the NeRF system already has an inherent drawback: it is prone to be overfitted to training views and struggles to grasp correct geometry when only sparse view inputs are available. Moreover, we experimentally find that blurred images lead to more severe overfitting in DeRF from sparse view because blur kernels introduce a more complex optimization process compared to standard NeRF. Due to the increased complexity, DeRF training suffers more structural distortion than general NeRFs when trained from sparse view, exhibiting further overfitting with floating artifacts as shown in our experiments. Although there are several works [4, 5, 6] to regularize radiance fields in sparse view scenarios, existing regularization methods are not effective in addressing the complex optimization issue of DeRF as demonstrated in comparative experiments using existing representative regularization techniques in sparse view NeRF and blur kernel of the DeRF, namely RegNeRF [6] and DP-NeRF [2]. Furthermore, in the DeRF system, it is challenging to use other data-deriven priors such as predicted depth supervision since available images do not ensure the confidence of the estimated values due to the inherent degradation of the given images. Therefore, our goal is to regularize the complex joint optimization of blur kernel and radiance fields for DeRF to enhance the structural and perceptual quality of radiance fields from sparse blurry images, overcoming the aforementioned challenging issues.
@@
In addition to geometric constraints, we propose the perceptual distillation (PD) as a perceptual prior to enhance the detailed texture of the radiance fields by taking advantage of the previously established image deblurring algorithm. Traditional image deblurring has shown significant performance improvements alongside the advancement of deep learning, demonstrating more enhanced details and textures. We believe that the sharp texture information from such deblurred images can be used as additional complementary information to achieve high fidelity in the Sparse-DeRF environment, where only a few degraded images are available to reconstruct the scene. However, while we can take the pre-filtered images with a pre-trained deep learning-based image deblurring model, the independence of image deblurring poses challenges in directly utilizing deblurred images as pixel-wise color supervision, due to inconsistency across the given images. This inconsistency comes from the inherent ill-posed property of the image deblurring that breaks the geometric and appearance consistency across the multi-view images of the single 3D scene. Hence, we impart the perceptual information of pre-filtered images to the radiance fields by distilling the features extracted from the deep learning-based image feature extractor. Extracted features enable the radiance fields to enhance perceptual quality by utilizing pre-deblurred textures.
@@ II-A Neural Rendering and Radiance Fields
II-B Radiance Fields in Practical Scenarios
@@
Another dominant area is constructing the NeRF from sparse view images, which is a practical environment considering real-world scenarios. Sparse view images incur the inherent drawback of neural networks in that the network is more likely to be overfitted to the given data distribution. This leads to inconsistent scene geometry in the mapped representations, typically manifested as incorrectly predicted structural information, such as elongated density artifacts in the rendered color and depth images from novel views. Several approaches have mitigated this issue involving additional prior knowledge or out-of-distribution data. InfoNeRF [40] adopts entropy minimization to probability density function (PDF) of density value along the ray density to make the shape of the PDF sharper. RegNeRF [6] utilizes the statistical depth smoothness of real-world geometry [7] on unobserved ray patches to reduce the artifact. Recently, FlipNeRF [5] considers flipped rays on the surface as supplement unseen rays to regularize the scene geometry. In other approaches, some works, such as PixelNeRF [41], and DietNeRF [42], exploit the semantic information extracted from deep image feature extractors to utilize the representative power of neural networks in feature level. FreeNeRF [4] tries to alleviate the overfitting problem based on an optimization perspective, imposing some restrictions on the frequency level.
In addition to sparse view settings, establishing NeRF from degraded images is recently emerging since the ideal training condition in images for NeRF often breaks in real-world scenarios. RawNeRF [43] denoises the internal noise of the camera sensor to construct high dynamic range (HDR) radiance fields from dark raw images and controls the camera exposure. Similarly, NaN [44] deals with burst noise in images, generating denoised images based on IBRNet [45], which is another image-based rendering approach. For more practical applications of NeRF in real-world, DeblurNeRF [1] firstly attempts to deal with two types of blur degradation in images, blur from camera motion and defocus, constructing deblurred neural radiance fields (DeRF) from only blurry images. They imitate the blurring process integrating the concept of blind deblurring in image deblurring with the NeRF system, modeling the blur kernel as pixel-wise independent ray transformation and composition weights to approximate the blurring process. Another representative approach is proposed by DP-NeRF[2], which imposes physical consistency across the images by modeling the blur kernel as the 3D rigid transformation of rays depending on each view, to approximate the actual blurring process in the camera more precisely. Recently several approaches [46, 47, 3] are also proposed in succession, attempting to improve the quality of the constructed DeRF. One of the most actively researched areas among those mentioned earlier is NeRF from blurry images, which often occurs when users take pictures with their own devices.
However, as we mentioned in Section I, the experimental setup of using only 20$\sim$30 blurry images, as in previous studies, is not practical. If we assume a scenario where users only have access to blurry images, it is more realistic to consider that only a few images are available for a specific scene and all of those images are blurry. Therefore, we propose a more practical scenario by combining DeRF and the sparse view setting, thereby enhancing real-world applicability.
@@ III-A Deblurred Neural Radiance Fields
where $\hat{B}_{c}$, $\hat{B}_{f}$, and $B$ are expected coarse, fine, and ground truth blurred color of the ray r, respectively.
The blur kernels are representatively modeled as a different type of transformation in each paper, [1, 2], which we will describe in the next paragraph. After the joint training of the approximated blur kernel and NeRF simultaneously, they can render the clean neural radiance fields, which we refer to as deblurred neural radiance fields (DeRF), by evaluating only NeRF thanks to the theoretical basis of blind deblurring.
Figure 1: Simple illustration for different blur kernel modeling of (a) Deblur-NeRF [1] and (b) DP-NeRF [2]. The main difference between the two kernels is the consistency between transformed rays induced from the blur kernel. Unlike Deblur-NeRF [1], which independently predict blurring rays for each pixel, DP-NeRF [2] predicts camera motion that makes the blur of each image and shares the motion across the entire pixels in the same image.
@@ III-B Blur Kernels in Deblurred Neural Radiance Fields
The core difference of the blur kernels between the two papers is a consideration of 3D consistency across the entire pixels in each image as we describe in Figure 1. Deblur-NeRF [1] designs the blur kernel as a deformable sparse kernel (DSK), which consists of the $n$ number of transformations depending on embedded latent features from each view and location of image pixels. The transformation is formulated as the 3D vector of ray origin and 2D vector of pixel coordinate, which is initialized within $N\times N$ window on the image plane as
@@
However, DP-NeRF [2] argues that the pixel-wise independent optimization of the blur kernel in [1] incurs inconsistency in 3D geometry and appearance. They utilize the physical intuition of actual blurred image acquisition in the camera process as an additional prior for the DeRF, to impose the constraints for constructing radiance fields while preserving 3D consistency. To directly model the actual camera motion as a 3D rigid transformation, they introduce scene-wise $SE(3)$ fields inspired by [29], [49] and Rodrigues’ formula [50]. Scene-wise rigid transformation of the camera is formulated by estimated screw axis $S_{s}\in\mathbb{R}^{6}$ through MLPs depending on only scene information as Eq.11.
@@
In this paper, we present a novel approach that includes geometric constraints and a perceptual prior through extensive experiments based on these two types of blur kernels from Ddeblur-NeRF [1] and DP-NeRF [2], which can be regarded as flexible and rigid kernels, respectively. The conducted experiments demonstrate the effectiveness of Sparse-DeRF on both kernels. Furthermore, the results also reveal a trade-off for the flexibility of the kernels, depending on the specific properties inherent in the diverse scenes.
@@ IV Sparse-DeRF
We find that it is not effective to construct the radiance fields based on naive NeRF, Deblur-NeRF [1], or DP-NeRF [2] from sparse view setting as we present in Section V-E. Although the DeRF usually recovers the high-frequency detail better than the naive NeRF in given views, it has become more fragmented in novel view synthesis, generating inaccurate scene geometry due to the complex joint optimization. The geometric error is represented as mapped RGB textures resembling painted walls in near or far depth and elongated density artifacts, which reveal challenges associated with accurate depth value prediction. However, existing representative regularization for the NeRF from sparse view [6] can not regularize complex joint optimization of the DeRF. We present experimental results in Section V-F1, which shows the difficulty of the previous regularization technique on the DeRF from sparse view. Hence, here we describe our method for regularizing optimization of the DeRF from sparse view to alleviate spatial ambiguity, which consists of two geometric constraints and a perceptual prior. Figure 2 illustrates the overall architecture of the Sparse-DeRF in detail, where (a),(b), and (c) of the Figure 2 indicate each main component of the Sparse-DeRF, respectively. Geometric constraints consist of surface smoothness (SS) and modulated gradient scaling (MGS) as we describe in Section IV-A and Section IV-B. A perceptual prior consists of a perceptual distillation (PD) which is described in Section IV-C.
Figure 2: Overall architecture of the Sparse-DeRF. Each color of integrated ray $\textbf{r}_{io}$ describes where the rays come from. $\hat{d}_{\Theta}(\textbf{r}_{i\sim i+k,j\sim j+k})$ indicates the rendered depth patch with $k\times k$ size. Note that, $k$ is set to $S_{ptc}$ as we described in the paper. The main component of the Sparse-DeRF consists of (a) surface smoothness (SS), (b) perceptual distillation (PD), and (c) modulated gradient scaling (MGS), which are indicated as gray, pink, and green colored components.
@@ IV-A Surface Smoothness on Integrated Unobserved Rays
In addition to leveraging previously unseen rays, we employ hidden rays derived from the characteristics of the estimated blur kernel, harnessing supplementary information exclusively presented in blurry inputs. The blur kernel, denoted as $h$ in Eq. 6, generates the ray transformation to approximate the color composition process of the blur, regardless of the kernel types, as demonstrated in Deblur-NeRF [1] and DP-NeRF [2]. Motivated by the commonality in the color composition process across various blur kernels, the kernel-induced transformed rays $\textbf{r}_{q}$ are introduced as additional unobserved rays, which are referred to as hidden rays, to enforce depth smoothness constraint. As the hidden rays $\textbf{r}_{q}$ are not directly presented in the training data for the DeRF system, they serve as supplementary out-of-distribution data for imposing depth smoothness, similar to the aforementioned unseen rays. In addition, the incorporation of depth smoothness on hidden rays effectively addresses geometric inconsistency across the 3D space of blurry images within the specified training view. This capability stems from the broader coverage of the estimated hidden rays, facilitated by the implicit inclusion of camera motion information in blurry images. Therefore, our integrated unobserved rays for depth smoothness are defined as an integrated set of unseen rays and hidden rays as

### v2

@@ Abstract
Recent studies construct deblurred neural radiance fields (DeRF) using dozens of blurry images, which are not practical scenarios if only a limited number of blurry images are available. This paper focuses on constructing DeRF from sparse-view for more pragmatic real-world scenarios. As observed in our experiments, establishing DeRF from sparse views proves to be a more challenging problem due to the inherent complexity arising from the simultaneous optimization of blur kernels and NeRF from sparse view. Sparse-DeRF successfully regularizes the complicated joint optimization, presenting alleviated overfitting artifacts and enhanced quality on radiance fields. The regularization consists of three key components: Surface smoothness, helps the model accurately predict the scene structure utilizing unseen and additional hidden rays derived from the blur kernel based on statistical tendencies of real-world; Modulated gradient scaling, helps the model adjust the amount of the backpropagated gradient according to the arrangements of scene objects; Perceptual distillation improves the perceptual quality by overcoming the ill-posed multi-view inconsistency of image deblurring and distilling the pre-deblurred information, compensating for the lack of clean information in blurry images. We demonstrate the effectiveness of the Sparse-DeRF with extensive quantitative and qualitative experimental results by training DeRF from 2-view, 4-view, and 6-view blurry images.
@@ I Introduction
In real-world scenarios, tackling the blurry images from camera motion is regarded to be important since users often encounter degraded images when capturing photos with their own devices due to the unintentional camera movement during exposure time. To solve this problem, several NeRF studies [1, 2, 3] have attempted to construct deblurred neural radiance fields (hereafter, DeRF) from blurry images using joint optimization of internal implicit blur kernel and radiance fields, but they use dozens of blurry images to train, which is actually not practical scenarios. The assumed experimental environments, where radiance fields are trained from about 20 to 30 blurry images, seem unlikely to occur in reality. Hence we delve into the practical consideration for situations where only blurry images are utilized. We reasoned that situations requiring the use of only blurry images would arise when the available images for reconstructing the desired 3D space are both very limited and blurry. Following this rationale, we propose a novel pragmatic scenario for radiance fields from blurry images that establish the DeRF from sparse view settings. Specifically, we set the 2, 4, and 6-view settings based on our consideration of the practical applications of research on generating radiance fields from blurry images.
Actually, the NeRF system already has an inherent drawback: it is prone to be overfitted to training views and struggles to grasp correct geometry when only sparse view inputs are available. Moreover, we experimentally find that blurred images lead to more severe overfitting in DeRF from sparse view because blur kernels introduce a more complex optimization process compared to standard NeRF. Due to the increased complexity, DeRF training suffers more structural distortion than general NeRFs when trained from sparse view, exhibiting further overfitting with floating artifacts as shown in our experiments. Although there are several works [4, 5, 6] to regularize radiance fields in sparse view scenarios, existing regularization methods are not effective in addressing the complex optimization issue of DeRF as demonstrated in comparative experiments using existing representative regularization techniques in sparse view NeRF and blur kernel of the DeRF, namely RegNeRF [6] and DP-NeRF [2]. Furthermore, in the DeRF system, it is challenging to use other data-deriven priors such as predicted depth supervision since available images do not ensure the confidence of the estimated values due to the inherent degradation of the given images. The inherent degradations cause inconsistency across multiple views, which hinder accurate geometry prediction in NeRF, when it is used as additional prior. This tendency becomes more pronounced as the number of input images increases. We validated the tendency through experimental results. Therefore, our goal is to regularize the complex joint optimization of blur kernel and radiance fields for DeRF to enhance the structural and perceptual quality of radiance fields from sparse blurry images, overcoming the aforementioned challenging issues without pretrained depth prior.
@@
In addition to geometric constraints, we introduce the perceptual distillation (PD) as a perceptual prior to enhance the detailed texture of the radiance fields by taking advantage of the established image deblurring algorithm. Traditional image deblurring has shown significant performance alongside the advancement of deep learning, demonstrating more enhanced details and textures. We believe that the sharp texture information from such deblurred images can be used as complementary information to achieve high fidelity in the Sparse-DeRF environment, where only a few degraded images are available to reconstruct the scene. However, while we can take the pre-deblurred images with a pre-trained deep learning-based image deblurring model, the independence of image deblurring poses challenges in directly utilizing deblurred images as pixel-wise color supervision, due to inconsistency across the given images. This inconsistency comes from the inherent ill-posed property of the image deblurring that breaks the geometric and appearance consistency across the multi-view images of the single 3D scene. Hence, we impart the perceptual information of pre-deblurred images to the radiance fields by distilling the features extracted from the deep learning-based image feature extractor. Extracted features enable the radiance fields to enhance perceptual quality by utilizing pre-deblurred textures.
@@ II-A Neural Rendering and Radiance Fields
II-B Radiance Fields from Degraded Images
Establishing NeRF from degraded images is recently emerging since the ideal training condition in images for NeRF often breaks in real-world scenarios. RawNeRF [37] denoises the internal noise of the camera sensor to construct high dynamic range (HDR) radiance fields from dark raw images and controls the camera exposure. Similarly, NaN [38] deals with burst noise in images, generating denoised images based on IBRNet [39], which is another image-based rendering approach. For more practical applications of NeRF in real-world, DeblurNeRF [1] firstly attempts to deal with two types of blur degradation in images, blur from camera motion and defocus, constructing deblurred neural radiance fields (DeRF) from only blurry images. They imitate the blurring process integrating the concept of blind deblurring in image deblurring with the NeRF system, modeling the blur kernel as pixel-wise independent ray transformation and composition weights to approximate the blurring process. Another representative approach is DP-NeRF [2], which is a key prior work that provided strong motivation for our current research and serves as the main baseline in our experiments. DP-NeRF [2] imposes physical consistency across the images by modeling the blur kernel as the 3D rigid transformation of rays depending on each view, to approximate the actual blurring process in the camera more precisely. Several approaches [40, 41, 3] are also proposed, attempting to improve the quality of the constructed DeRF with simple motion interpolation and progressive blur kernel estimation. With the emergence of the Gaussian splatting [42], researches such as [43, 44, 45] have also been published in this field, demonstrating that research remains active and ongoing.
One of the most actively researched areas among those mentioned earlier is NeRF from blurry images, which often occurs when users take pictures with their own devices. However, as we mentioned in Section I, the experimental setup of using only 20$\sim$30 blurry images, as in previous studies, is not practical. If we assume a scenario where users only have access to blurry images, it is more realistic to consider that only a few images are available for a specific scene and all of those images are blurry. Therefore, we propose a more practical scenario by combining DeRF and the sparse view setting, thereby enhancing real-world applicability.
II-C Radiance Fields from Sparse View
@@
Another dominant area is constructing the NeRF from sparse view images, which is a practical environment considering real-world scenarios. Sparse view images incur the inherent drawback of neural networks in that the network is more likely to be overfitted to the given data distribution. This leads to lack of the information for dense scene geometry in the mapped representations, typically manifested as incorrectly predicted structure, such as elongated density artifacts in the rendered color and depth images from novel views. Several approaches have mitigated this issue involving additional prior knowledge or out-of-distribution data. InfoNeRF [49] adopts entropy minimization to probability density function (PDF) of density value along the ray density to make the shape of the PDF sharper. RegNeRF [6] utilizes the statistical depth smoothness of real-world geometry [7] on unobserved ray patches to reduce the artifact. Recently, FlipNeRF [5] considers flipped rays on the surface as supplement unseen rays to regularize the scene geometry. In other approaches, some works, such as PixelNeRF [50], and DietNeRF [51], exploit the semantic information extracted from deep image feature extractors to utilize the representative power of neural networks in feature level. FreeNeRF [4] tries to alleviate the overfitting problem based on an optimization perspective, imposing some restrictions on the frequency level.
There have been various studies attempting to address the challenge of accurate geometry prediction in sparse-view settings by utilizing depth priors obtained from pretrained depth estimation networks or point clouds generated by structure-from-motion (SfM). For example, SparseNeRF [52] tackled this issue by distilling depth ranking information, while SPARF [53] incorporated geometric warping to refine the depth priors for unseen views. DSNeRF [54] leveraged point clouds obtained from SfM as depth priors and Dense-Depth-Prior [55] further employed depth completion techniques to reduce the sparsity of the SfM point cloud, resulting in denser depth priors. StructNeRF [56] incorporate superpixel segmentation to refine the depth priors even more effectively. While these approaches attempt to address the lack of information by using depth priors, our proposed practical experimental settings that use the sparse blurry images reveal significant limitations. As mentioned earlier, the presence of blur in each image makes predicted depth inaccurate and the inconsistent across the multiple views, which hinders their usability as a prior.
@@ III-A Deblurred Neural Radiance Fields
where $\hat{B}_{c}$, $\hat{B}_{f}$, and $B$ are expected coarse, fine, and ground truth blurred color of the ray r, respectively. The blur kernels are representatively modeled as a different type of transformation in each paper, [1, 2].
Figure 1: Simple illustration for different blur kernel modeling of (a) Deblur-NeRF [1] and (b) DP-NeRF [2]. The main difference between the two kernels is the consistency between transformed rays induced from the blur kernel.
@@ III-B Blur Kernels in Deblurred Neural Radiance Fields
The core difference of the blur kernels between the two papers [1, 2] is a consideration of 3D consistency across the entire pixels in each image as we illustrate in Figure 1. Deblur-NeRF [1] designs the blur kernel as a deformable sparse kernel (DSK), which consists of the $n$ number of transformations depending on embedded latent features from each view and location of image pixels. The transformation is formulated as the 3D vector of ray origin and 2D vector of pixel coordinate, which is initialized within $N\times N$ window on the image plane as
@@
However, DP-NeRF [2] argues that the pixel-wise independent optimization of the blur kernel in [1] incurs inconsistency in 3D geometry and appearance. They utilize the physical intuition of actual blurred image acquisition in the camera process as an additional prior for the DeRF, to impose the constraints for constructing radiance fields while preserving 3D consistency. To ensure 3D consistency, unlike Deblur-NeRF [1], they made the blur process shared within a single image and defined it as a learnable set of camera motions. To directly model the actual camera motion as a 3D rigid transformation, they introduce scene-wise $SE(3)$ fields inspired by [29], [58] and Rodrigues’ formula [59]. Scene-wise rigid transformation of the camera is formulated by estimated screw axis $S_{s}\in\mathbb{R}^{6}$ through MLPs depending on only scene information as Eq.11.
@@
In this paper, we propose a novel approach that consists of geometric constraints and a perceptual prior through extensive experiments with these two types of blur kernels from Deblur-NeRF [1] and DP-NeRF [2]. The conducted experiments demonstrate the effectiveness of Sparse-DeRF on both kernels.
@@ IV Sparse-DeRF
We find that it is not effective to construct the radiance fields based on naive NeRF, Deblur-NeRF [1], or DP-NeRF [2] from sparse view setting as we present in Section V-E. Although the DeRF usually recovers the high-frequency detail better than the naive NeRF in given views, it has become more fragmented in novel view synthesis, generating inaccurate scene geometry due to the complex joint optimization and overfitting. The geometric error is represented as mapped RGB textures resembling painted walls in near or far depth and elongated density artifacts, which reveal challenges associated with accurate depth value prediction. However, existing representative regularization for the NeRF from sparse view [6] can not regularize complex joint optimization of the DeRF. We present experimental results in Section V-F1, which shows the difficulty of the previous regularization technique on the DeRF from sparse view. Hence, here we describe our method for regularizing optimization of the DeRF from sparse view to alleviate spatial ambiguity, which consists of two geometric constraints and a perceptual prior. Figure 2 illustrates the overall architecture of the Sparse-DeRF in detail, where (a),(b), and (c) of the Figure 2 indicate each main component of the Sparse-DeRF, respectively. Geometric constraints consist of surface smoothness (SS) and modulated gradient scaling (MGS), and a perceptual prior consists of a perceptual distillation (PD).
Figure 2: Overall architecture of the Sparse-DeRF. The main component of Spare-DeRF consists of three components, Surface Smoothness (SS), Modulated Gradient Scaling (MGS), and Perceptual Distillation (PD), which are denoted as (a), (b), (c) in the figure. Note that, NeRF network is shared to predict the color of each sampled points along the ray in hidden rays, integrated unobserved rays and rays for patch rendering.
@@ IV-A Surface Smoothness on Integrated Unobserved Rays
In addition to leveraging previously unseen rays, we employ hidden rays derived from the characteristics of the estimated blur kernel, harnessing supplementary information exclusively presented in blurry inputs. The blur kernel, denoted as $h$ in Eq. 6, generates the ray transformation to approximate the color composition process of the blur, regardless of the kernel types, as demonstrated in Deblur-NeRF [1] and DP-NeRF [2]. Motivated by the commonality in the color composition process across various blur kernels, these kernel-induced transformed rays $\textbf{r}_{q}$ are introduced as additional unobserved rays, which are referred to as hidden rays, to enforce depth smoothness constraint to the blur kernel. As the hidden rays $\textbf{r}_{q}$ are not directly presented in the training data for the DeRF system, they serve as supplementary data for imposing depth smoothness regularization effect to the blur kernel and stabilize the complex optimization of blur kernel and radiance fields. In addition, the incorporation of depth smoothness on hidden rays effectively help alleviate the geometric inconsistency across the 3D space of blurry images within the specified training view. Therefore, our integrated unobserved rays for depth smoothness are defined as an integrated set of unseen rays and hidden rays as